sin(2n+1)A.sinA = sin^2(n+1)A-sin^2(nA)

Grade:11

1 Answers

Praneetha reddy
11 years ago

By using this formula,

sin2A - sin2B = sin ( A + B ) sin ( A - B )

RHS = sin2(nA +A) - sin2(nA) = sin( nA + A +nA ) sin ( nA + A - nA) = sin (2n + 1) A sin A = LHS